# GM-MoE: Low-Light Enhancement with Gated-Mechanism Mixture-of-Experts

### v2

@@ Abstract
Low-light enhancement has wide applications in autonomous driving, 3D reconstruction, remote sensing, surveillance, and so on, which can significantly improve information utilization. However, most existing methods lack generalization and are limited to specific tasks such as image recovery. To address these issues, we propose Gated-Mechanism Mixture-of-Experts (GM-MoE), the first framework to introduce a mixture-of-experts network for low-light image enhancement. GM-MoE comprises a dynamic gated weight conditioning network and three sub-expert networks, each specializing in a distinct enhancement task. Combining a self-designed gated mechanism that dynamically adjusts the weights of the sub-expert networks for different data domains. Additionally, we integrate local and global feature fusion within sub-expert networks to enhance image quality by capturing multi-scale features. Experimental results demonstrate that the GM-MoE achieves superior generalization with respect to 25 compared approaches, reaching state-of-the-art performance on PSNR on 5 benchmarks and SSIM on 4 benchmarks, respectively.
Keywords Low-Light Image Enhancement  $\cdot$ Multi-Expert System  $\cdot$ Gated Mechanism
@@ 1 Introduction
Figure 1: Given a low-light image, our GE-MoE achieves better performance (for both object and whole scene) compared with LightenDiffusion [15]
@@
To address these issues, we propose an innovative Gated-Mechannism Mixture-of-Experts (GM-MoE) system for low-light image enhancement. The method is based on an improved U-Net [18] architecture, incorporating a gated mechanism expert network with dynamic weight adjustment to adapt to photo inputs from different data domains. GM-MoE consists of three sub-expert networks, each of which is used to solve different image enhancement tasks, namely color correction, detail recovery, and problems. The gating mechanism assigns appropriate weights to each sub-expert based on the different lighting and scene conditions of the image, achieving a balance between the image enhancement problems and thus achieving the best image enhancement under different lighting conditions and scenes.
As shown in Fig. 2, our framework achieves a higher PSNR and SSIM in multiple data sets compared to other state-of-the-art methods discussed in the literature.
Figure 2: The comparison results among GM-MoE and the SOTA low-light image enhancement methods on the LOL-v1 , LOLv2-Real , LOLv2-Synthetic and LSRW-Huawei benchmarks. GM-MoE outperforms all of compared approaches on both PSNR and SSIM metrics.
The main contributions of GM-MoE are summarized as follows:
  • First application of GM-MoE to low-light image enhancement. Our method integrates a dynamically gated weight adjustment network with a multi-expert module, enabling effective generalization across different data domains.
  • Dynamic gate-controlled weight adjustment. We design a novel mechanism that allows the MoE framework to dynamically adapt to varying lighting conditions, optimizing enhancement results.
  • Superior performance across multiple datasets and downstream tasks. Extensive experiments demonstrate that GM-MoE surpasses SOTA methods on different metrics across five datasets while maintaining strong generalization ability.
Figure 3: Overview of the proposed GM-MOE. (a) The GM-MoE module comprises a dynamically gated weight adjustment network and three specialized sub-expert networks. (b) The overall network adopts a U-Net-like encoder-decoder architecture. Given an input image, it first undergoes processing through the Shallow Feature Extraction Block (SFEB). Then, the GM-MoE module facilitates multi-scale feature fusion via multiple downsampling and upsampling operations, ultimately generating the enhanced output image.
@@ 2.2 Multi-Expert Systems
In the field of low-light image enhancement, there is a lack of MoE applications. At the same time, the fact that the expert network module can solve different problems simultaneously makes it very advantageous in low-light image enhancement. Based on this, in order to solve the current problems of low-light enhancement models, i.e., most of the models can only deal with a single problem of low-light image restoration but cannot address multiple problems faced by low-light images simultaneously, nor can they be adjusted according to the different lighting conditions of the input photos and the corresponding processing approach.
As shown in Fig. 3, We design a dynamic weight-adjusted multi-expert network based on the U-Net architecture, with each sub-expert module targeting the problems of color distortion, detail loss, and low contrast. To solve the problem of poor generalization ability of the existing models, we incorporate a dynamic weight adjustment network that dynamically adjusts the weights of the three sub-networks according to the input photographs to adapt to photographs from different data domains.
@@ 3 Method
In the model we designed, the overall architecture is as follows: The input dark image $I\in\mathbb{R}^{H\times W\times 3}$ is first processed by SFEB to obtain low-level features $X_{0}\in\mathbb{R}^{H\times W\times C}$, where $H$ and $W$ denote the spatial dimensions of the image and $C$ denotes the number of channels. These features are then input into a module similar to the U-Net [18] network architecture, which contains encoder and decoder layers to further extract deeper features $F_{d}\in\mathbb{R}^{H\times W\times 2C}$. The GM-MOE module is introduced in the encoder and decoder of each layer. During encoding, the encoder of each layer compresses the image features by gradually reducing the spatial dimension and increasing the channel capacity, while the decoder gradually restores the image resolution by upsampling the low-resolution feature map $F_{l}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 8C}$ to progressively recover the image resolution. To optimize feature preservation and recovery, pixel-unshuffle and pixel-shuffle techniques are introduced to improve the effects of downsampling and upsampling, respectively. To aid low-light image feature recovery, the initial features are preserved and fused between the encoder and decoder via skip connections. Each layer of the GM-MOE module is responsible for fusing the lower-level features of the encoder with the higher-level features of the decoder, thereby enriching the structure and texture details of the image. In the final stage, the deep features $F_{d}$ further enhances the detailed features at the spatial resolution, and the residual image generated by the convolution operation $R$ is added to the input image $I$ to obtain the final enhanced image $\hat{I}=I+R$.
@@ 3.1 Gated-Mechanism Mixture-of-Experts
As shown in Fig. 3, the color restoration expert network (Expert1, also named Net1) is used to restore the color information of images under low light conditions. Our designed color restoration subnetwork first learns the image information while reducing the dimension through convolution and downsampling operations. Then, the extracted features are restored to the image resolution through max pooling and upsampling, and bilinear interpolation is used to ensure natural image transition during restoration:
@@
As shown in Fig. 3, the advanced feature enhancement subnetwork (Expert3, also named Net3) improves image quality through convolution, multi-scale feature extraction, a gating mechanism, and an attention mechanism. The input image is passed through a multi-scale convolution to extract and fuse features. These fused features are then processed further by a gating network (SG) and a channel attention mechanism (SCA). Finally, the enhanced features are added back to the input image via a residual connection to preserve the original details. This method can effectively adapt to low-light scenes and improve image quality by dynamically adjusting the weights of the expert network.
@@ 3.2 Shallow Feature Extraction Module
To improve the effectiveness of image feature extraction and suppress invalid features, we design a multi-scale feature enhancement module (SFEB, Shallow Feature Enhancement Block) to process the input feature map $X\in\mathbb{R}^{C\times H\times W}$. The SFEB generates a feature map $F_{1}\in\mathbb{R}^{C\times H\times W}$ through a $3\times 3$ depth separable convolution, as shown in Fig.  4 . In addition, SFEB also obtains a feature map $F_{2}\in\mathbb{R}^{C\times H\times W}$ using different sizes of convolution kernels through hole convolution (dilation rates) to capture multi-scale spatial information.
SFEB uses two $1\times 1$ convolutions to compress the channel numbers of $F_{1}$ and $F_{2}$ to form the fused feature map $F_{e}\in\mathbb{R}^{C^{\prime}\times H\times W}$, which fuses different feature information.
To introduce the attention mechanism, SFEB performs global pooling on the fused feature map $F_{e}$ to obtain the channel-weighted features $A_{avg}$ and $A_{max}$. Then, the attention map $A\in\mathbb{R}^{C^{\prime}\times H\times W}$ is generated by channel concatenation and a $7\times 7$ convolution to enhance the features of key regions:
$F_{w}=F_{1}^{{}^{\prime}}\odot{A}_{1}+F_{2}^{{}^{\prime}}\odot{A}_{2}$ (5)
Finally, the output feature map $Y$ is obtained by element-wise multiplication of the input feature $X$ and the attention map ${A}^{{}^{\prime}}$:
${Y}={X}\odot{A}^{{}^{\prime}}$ (6)
@@ 4.1 Datasets and Implementation Details
PairLIE [60] 23.526 0.755 19.885 0.778 19.074 0.794 0.33
LLFormer [61] 25.758 0.823 20.056 0.792 24.038 0.909 24.55
QuadPrior [62] 22.849 0.800 20.592 0.811 16.108 0.758 1252.75
@@
SID [46] 17.47 0.6518 16.33 0.6127
RF [47] 19.05 0.6374 18.77 0.6302
UFormer [48] 19.77 0.6432 19.77 0.6432
EnGAN [49] 20.22 0.7012 20.71 0.6588
Restormer [16] 22.61 0.7254 21.20 0.6774
@@
LIME [29] 17.00 0.3816 13.53 0.3321
MF [52] 18.26 0.4283 15.44 0.3997
NPE [53] 17.08 0.3905 14.86 0.3738
SRIE [54] 13.42 0.4282 13.26 0.1396
RetinexNet [37] 19.98 0.6879 19.86 0.6501
Kind [38] 16.58 0.5690 11.52 0.3827
Kind++ [55] 15.43 0.5695 14.79 0.4749
MIRNet [56] 19.98 0.6085 17.10 0.5022
SNR-Net [57] 20.67 0.5910 17.54 0.4822
Bread [58] 19.20 0.6179 14.70 0.4867
LightenDiffusion [15] 18.555 0.539 - -
@@
Implementation Details: The GM-MoE was developed using the PyTorch framework and trained on an NVIDIA 4090 GPU over a period of three days. The training process began with an initial learning rate of $1.0\times 10^{-3}$, which was managed using a multi-step scheduler. The Adam [64] optimizer, configured with a momentum parameter of 0.9, was used for optimization. During training, input images were resized to $256\times 256$ pixels and subjected to data augmentation techniques, including random rotations and flips, to enhance model generalization. A batch size of 4 was maintained, and the training regimen consisted of a total of $2.0\times 10^{6}$ iterations. Performance was evaluated using Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) [65] as the primary metrics.
@@ 4.2 Low-light Image Enhancement
Quantitative Results. Among the deep learning methods for low-light image enhancement are SID [46], RF [47], UFormer [48], EnGAN [49], Restormer [16], Retinexformer [50], DeepUPE [51], RetinexNet [37], Kind [38], Kind++ [55], MIRNet [56], SNR-Net [57], Bread [58], DPEC [59], as well as traditional methods such as, LIME [29], MF [52], NPE [53], and SRIE [54], GM-MoE achieved an overall performance improvement on all datasets. Tab. 2 shows the quantitative comparison results of GM-MoE with a variety of SOTA image enhancement algorithms. GM-MoE achieved PSNR improvements of 1.5, 1.24, and 0.11 dB on the LOL-v1 [37], LOLv2-Real [45], and LOLv2-Synthetic [45] datasets, respectively, compared to the second-ranked model in each dataset, significantly improving image quality. This showed that GM-MoE performed well on these classic datasets, consistently outperforming other methods in terms of detail recovery and color enhancement, both in synthetic and real low-light conditions. In addition, as shown in Tab. 2, GM-MoE also achieved significant improvements over other SOTA methods on the LSRW-Huawei [63] and LSRW-Nikon [63] datasets, achieving PSNR improvements of 0.94 dB and 1.42 dB over the next best method, Restormer [16], respectively. These two datasets contained a large amount of noise and artifacts, and GM-MoE was able to effectively reduce artifacts and recover image details and features in high-noise environments. For more results please refer to supplementary material.
@@
Is the Effectiveness Among Multiple Experts and the Performance Complementary? To explore the role of each expert module (Expert1,Expert2, Expert3) and its complementary role in image enhancement, we gradually added each subnetwork to the model and performed ablation experiments. After adding Expert1, compared with Experiment 2, PSNR increased by 1.08 dB (LOL-v2-real), respectively. After adding Expert2, compared with Experiment 3, PSNR increased by 0.76 dB (LOL-v2-real) and 0.79 dB (LOL-v2-syn), and SSIM increased by 0.0575 and 0.0891, respectively. Subsequently, we removed Expert1, Expert2, and Expert3 respectively through ablation experiments (Experiments 5–7), and a decrease in performance was observed in all cases. This shows that the expert modules can work together in synergy after being integrated into a module to solve the image enhancement problems they are designed for, improving the overall image restoration effect.
Does the Dynamic Gating Weight Adjustment Network Improve Generalization Ability? In the complete model, we also added a dynamic gating weight adjustment network to adaptively assign weights to each subnetwork. Compared with the model without dynamic weight adjustment, incorporating this mechanism can effectively dynamically adjust each expert’s information according to photos of different data domains, improving the generalization ability of the model across different datasets.
ID Baseline SFEB Net1 Net2 Net3 WA LOL-v2-real [45] LOL-v2-syn [45]
PSNR SSIM PSNR SSIM
1 19.45 0.7079 20.35 0.7431
2 20.27 0.7236 23.44 0.7646
3 21.35 0.7446 24.35 0.8436
4 22.11 0.8021 25.14 0.9327
5 23.23 0.8045 26.08 0.9351
6 23.31 0.8054 26.12 0.9362
7 23.35 0.8055 26.15 0.9366
8 23.65 0.8060 26.29 0.9371
@@
Qualitative results.
As shown in the ablation study results in Fig. 8, images generated using only the baseline model exhibit noticeable blurring and loss of detail. The sequential integration of SFEB, Net1, Net2, Net3, and the dynamic gate-weighted adjustment module progressively enhances the model’s ability to recover low-light images. Each component contributes uniquely to the overall performance, and the complete model achieves the best results.

### v5

@@ Abstract
Low-light enhancement has wide applications in autonomous driving, 3D reconstruction, remote sensing, surveillance, and so on, which can significantly improve information utilization. However, most existing methods lack generalization and are limited to specific tasks such as image recovery. To address these issues, we propose Gated-Mechanism Mixture-of-Experts (GM-MoE), the first framework to introduce a mixture-of-experts network for low-light image enhancement. GM-MoE comprises a dynamic gated weight conditioning network and three sub-expert networks, each specializing in a distinct enhancement task. Combining a self-designed gated mechanism that dynamically adjusts the weights of the sub-expert networks for different data domains. Additionally, we integrate local and global feature fusion within sub-expert networks to enhance image quality by capturing multi-scale features. Experimental results demonstrate that the GM-MoE achieves superior generalization compared to over 20 existing approaches, reaching state-of-the-art performance on PSNR on 5 benchmarks and SSIM on 4 benchmarks, respectively. Code is available at: https://github.com/Sameenok/gm-moe-lowlight-enhancement.git
@@ 1 Introduction
Figure 1: Given a low-light image, our GM-MoE achieves better performance (for both object and whole scene) compared with LightenDiffusion [20].
@@
To address these issues, we propose an innovative Gated-Mechannism Mixture-of-Experts (GM-MoE) system for low-light image enhancement. The method is based on an improved U-Net [41] architecture, incorporating a gated mechanism expert network with dynamic weight adjustment to adapt to photo inputs from different data domains. GM-MoE consists of three sub-expert networks, each of which is used to solve different image enhancement tasks, namely color correction, detail recovery, and problems. The gating mechanism assigns appropriate weights to each sub-expert based on the different lighting and scene conditions of the image, achieving a balance between the image enhancement problems and thus achieving the best image enhancement under different lighting conditions and scenes. As shown in Fig. LABEL:comparison1, our framework achieves a higher PSNR and SSIM in multiple data sets compared to other state-of-the-art methods discussed in the literature.
Figure 2: The comparison results among GM-MoE and the SOTA low-light image enhancement methods on the LOL-v1 , LOLv2-Synthetic and LSRW-Huawei benchmarks. GM-MoE outperforms all of compared approaches on both PSNR and SSIM metrics.
In summary, our contributions are as follows:
  • We are the first to apply GM-MoE to low-light image enhancement, proposing a method that combines a dynamic gating weight adjustment network with a multi-expert module to achieve effective generalization across different data domains.
  • We propose a dynamic gating mechanism that adaptively adjusts the MoE weights according to varying lighting conditions, thereby optimizing image enhancement.
  • Our model achieves superior performance across multiple datasets and downstream tasks. Extensive experiments demonstrate that GM-MoE surpasses SOTA methods on different metrics across five datasets while maintaining strong generalization ability.
Figure 3: Overview of the proposed GM-MOE. (a) The GM-MoE module comprises a gated weight generation network and three specialized sub-expert networks. (b) The overall network adopts a U-Net-like encoder-decoder architecture. Given an input image, it first undergoes processing through the Shallow Feature Extraction Block (SFEB). Then, the GM-MoE module facilitates multi-scale feature fusion via multiple downsampling and upsampling operations, ultimately generating the enhanced output image.
@@ 2.2 Multi-Expert Systems
In the field of low-light image enhancement, there is a lack of MoE applications. The expert network module possesses the capability to simultaneously address multiple low-light image enhancement challenges, a characteristic that grants it unique advantages in the field of low-light image enhancement. Moreover, existing methods exhibit significant limitations: most low-light enhancement models can only achieve satisfactory restoration results for a single problem and lack the ability to adaptively adjust processing strategies based on the illumination conditions of the input image.
To address the limitations of existing methods and leverage the unique strengths of expert networks, this paper proposes a novel Multi-Expert Low-Light Enhancement Network. As illustrated in Fig. 3, we design a dynamically weighted multi-expert network based on the U-Net architecture, where three specialized sub-expert modules are optimized for color distortion, detail loss, and low-contrast issues, respectively. To overcome the insufficient generalization ability of existing models, we innovatively introduce a Dynamic Weight Adjustment Network, which automatically adjusts the weight allocation of the three sub-networks based on input image features, thereby achieving adaptive enhancement for images from different data domains.
@@ 3 Method
In the model we designed, the overall architecture is as follows: The input dark image $I\in\mathbb{R}^{H\times W\times 3}$ is first processed by SFEB to obtain low-level features $X_{0}\in\mathbb{R}^{H\times W\times C}$, where $H$ and $W$ denote the spatial dimensions of the image and $C$ denotes the number of channels. These features are then input into a module similar to the U-Net [41] network architecture, which contains encoder layers to further extract deeper features $F_{d}\in\mathbb{R}^{H\times W\times 2C}$. The GM-MOE module is introduced in the encoder and decoder of each layer. During encoding, the encoder of each layer compresses the image features by gradually reducing the spatial dimension and increasing the channel capacity, while the decoder gradually restores the image resolution by upsampling the low-resolution feature map $F_{l}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 8C}$ to progressively recover the image resolution. To optimize feature recovery, pixel-shuffle techniques are introduced to improve the effects of upsampling. To aid low-light image feature recovery, the initial features are preserved and fused between the encoder and decoder via skip connections. Each layer of the GM-MOE module is responsible for fusing the lower-level features of the encoder with the higher-level features of the decoder, thereby enriching the structure and texture details of the image. In the final stage, the deep features $F_{d}$ further enhances the detailed features at the spatial resolution, and the residual image generated by the convolution operation $R$ is added to the input image $I$ to obtain the final enhanced image $\hat{I}=I+R$.
@@ 3.1 Gated-Mechanism Mixture-of-Experts
As shown in Figure 3, the color restoration expert network (Expert1, also named Net1) is used to restore the color information of images under low light conditions. The color restoration subnetwork we designed employs pooling operations to focus on key color features during downsampling while simultaneously learning image information. Additionally, deconvolution operations are utilized to precisely restore image details, with nonlinear interpolation adopted to ensure smooth and natural color transitions:
@@
As shown in Fig. 3, the advanced feature enhancement subnetwork (Expert3, also named Net3) improves image quality through convolution, multi-scale feature extraction, a gating mechanism, and an attention mechanism. The input image is passed through a multi-scale convolution to extract and fuse features. These fused features are then processed further by a gating network (SG) and a channel attention mechanism (SCA). Finally, the enhanced features are added back to the input image via a residual connection to preserve the original details.
This method can effectively adapt to low-light scenes and improve image quality by dynamically adjusting the weights of the expert network.
@@ 3.2 Shallow Feature Extraction Module
To improve the effectiveness of image feature extraction and suppress invalid features, we design a multi-scale feature enhancement module (SFEB, Shallow Feature Enhancement Block) to process the input feature map $X\in\mathbb{R}^{C\times H\times W}$. The SFEB generates a feature map $F_{1}\in\mathbb{R}^{C\times H\times W}$ through a $3\times 3$ depth separable convolution, as shown in Fig. 4 . In addition, SFEB also obtains a feature map $F_{2}\in\mathbb{R}^{C\times H\times W}$ using different sizes of convolution kernels through hole convolution (dilation rates) to capture multi-scale spatial information.
SFEB uses two convolutions to compress the channel numbers of $F_{1}$ and $F_{2}$ to form the fused feature map $F_{e}\in\mathbb{R}^{C^{\prime}\times H\times W}$, which fuses different feature information.
To introduce the attention mechanism, SFEB performs global pooling on the fused feature map $F_{e}$ to obtain the channel-weighted features $A_{avg}$ and $A_{max}$. Then, the attention map $F_{w}\in\mathbb{R}^{C^{\prime}\times H\times W}$ is generated by channel concatenation and a $7\times 7$ convolution to enhance the features of key regions:
$F_{w}=F_{1}^{{}^{\prime}}\odot{A}_{avg}+F_{2}^{{}^{\prime}}\odot{A}_{max}$ (5)
Finally, the output feature map $Y$ is obtained by element-wise multiplication of the input feature $X$ and the attention map $F_{w}$:
${Y}={X}\odot F_{w}$ (6)
@@ 4.1 Datasets and Implementation Details
PairLIE [9] 23.53 0.755 19.89 0.778 19.07 0.794 0.33
LLFormer [67] 25.76 0.823 20.06 0.792 24.04 0.909 24.55
QuadPrior [49] 22.85 0.800 20.59 0.811 16.11 0.758 1252.75
3DLUT[66] 14.35 0.445 17.59 0.721 18.04 0.800 0.59
Sparse[62] 17.20 0.640 20.06 0.816 22.05 0.905 1.08
RUAS[40] 18.23 0.720 18.37 0.723 16.55 0.652 1.03
DRBN [61] 20.13 0.830 20.29 0.831 23.22 0.927 1.83
@@
SID [3] 17.47 0.652 16.33 0.613
RF [25] 19.05 0.637 18.77 0.630
UFormer [50] 19.77 0.643 19.77 0.643
EnGAN [22] 20.22 0.701 20.71 0.659
Restormer [65] 22.61 0.725 21.20 0.677
@@
LIME [12] 17.00 0.382 13.53 0.332
MF [7] 18.26 0.428 15.44 0.400
NPE [48] 17.08 0.391 14.86 0.374
SRIE [8] 13.42 0.428 13.26 0.140
RetinexNet [52] 19.98 0.688 19.86 0.650
Kind [68] 16.58 0.569 11.52 0.383
Kind++ [69] 15.43 0.570 14.79 0.475
MIRNet [64] 19.98 0.609 17.10 0.502
SNR-Net [54] 20.67 0.591 17.54 0.482
Bread [11] 19.20 0.618 14.70 0.487
LightenDiffusion [20] 18.56 0.539 - -
3DLUT[66] 18.12 0.659 17.81 0.629
Sparse[62] 20.33 0.699 20.19 0.657
RUAS[40] 20.46 0.704 20.88 0.664
DRBN [61] 20.61 0.710 21.07 0.670
@@
Implementation Details: The GM-MoE was developed using the PyTorch framework and trained on an NVIDIA 4090 GPU. The training process began with an initial learning rate of $1.0\times 10^{-3}$, which was managed using a multi-step scheduler. The Adam [35] optimizer, configured with a momentum parameter of 0.9, was used for optimization. During training, input images were resized to $256\times 256$ pixels and subjected to data augmentation techniques, including random rotations and flips, to enhance model generalization. A batch size of 4 was maintained, and the training regimen consisted of a total of $2.0\times 10^{6}$ iterations. Performance was evaluated using Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) [51] as the primary metrics.
@@ 4.2 Low-light Image Enhancement
Quantitative Results. Among the deep learning methods for low-light image enhancement are SID [3], RF [25], UFormer [50], EnGAN [22], Restormer [65], Retinexformer [2], DeepUPE [46], RetinexNet [52], Kind [68], Kind++ [69], MIRNet [64], SNR-Net [54], Bread [11], DPEC [47], as well as traditional methods such as, LIME [12], MF [7], NPE [48], and SRIE [8], GM-MoE achieved an overall performance improvement on all datasets. Tab. 1 shows the quantitative comparison results of GM-MoE with a variety of SOTA image enhancement algorithms. GM-MoE achieved PSNR improvements of 0.9, 0.76, and 0.11 dB on the LOL-v1 [52], LOLv2-Real [62], and LOLv2-Synthetic [62] datasets, respectively, compared to the second-ranked model in each dataset, significantly improving image quality. This showed that GM-MoE performed well on these classic datasets, consistently outperforming other methods in terms of detail recovery and color enhancement, both in synthetic and real low-light conditions. In addition, as shown in Tab. 2, GM-MoE also achieved significant improvements over other SOTA methods on the LSRW-Huawei [14] and LSRW-Nikon [14] datasets, achieving PSNR improvements of 0.94 dB and 1.42 dB over the next best method, Restormer [65], respectively. These two datasets contained a large amount of noise and artifacts, and GM-MoE was able to effectively reduce artifacts and recover image details and features in high-noise environments. For more results please refer to supplementary material.
@@
Is the Effectiveness Among Multiple Experts and the Performance Complementary? To explore the role of each expert module (Expert1,Expert2, Expert3) and its complementary role in image enhancement, we gradually added each subnetwork to the model and performed ablation experiments. After adding Expert1, compared with Experiment 2, PSNR increased by 1.08 dB (LOL-v2-real) . After adding Expert2, compared with Experiment 3, PSNR increased by 0.76 dB (LOL-v2-real) and 0.79 dB (LOL-v2-syn), and SSIM increased by 0.0575 and 0.0891, respectively. Subsequently, we removed Expert1, Expert2, and Expert3 respectively through ablation experiments (Experiments 5–7), and a decrease in performance was observed in all cases. This shows that the expert modules can work together in synergy after being integrated into a module to solve the image enhancement problems they are designed for, improving the overall image restoration effect.
Does the Gated Weight Generation Network Improve Generalization Ability? In the complete model, we further introduce a gated weight generation network. Compared to models without this mechanism, it dynamically adjusts the weights of individual experts based on images from different data domains, thereby enhancing cross-domain generalization capability.
ID Baseline SFEB Net1 Net2 Net3 GM LOL-v2-real [62] LOL-v2-syn [62]
PSNR
SSIM
PSNR
SSIM
1
19.45
0.7079
20.35
0.7431
2
20.27
0.7236
23.44
0.7646
3
21.35
0.7446
24.35
0.8436
4
22.11
0.8021
25.14
0.9327
5
23.23
0.8045
26.08
0.9351
6
23.31
0.8054
26.12
0.9362
7
23.35
0.8055
26.15
0.9366
8
23.65
0.8060
26.29
0.9371
@@
Qualitative results. As shown in the ablation study results in Fig. 8, images generated using only the baseline model exhibit noticeable blurring and loss of detail. The sequential integration of SFEB, Net1, Net2, Net3, and the gated weight generation module (GM) progressively enhances the model’s ability to recover low-light images. Each component contributes uniquely to the overall performance, and the complete model achieves the best results.